# A characterization of Fano type varieties

### 08 Mar 2026 v1

@@ 1. Introduction
In [zhuang2024direct], Zhuang proved that pure images of klt type singularities are of klt type. The following klt type criterion is crucial is his proof.
@@ Acknowledgments
The author acknowledges Professor Mao Sheng for the financial support through the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@
Let $f:X\to Y$ be a projective morphism with connected fibers between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. There are effective $\mathbb{Q}$-divisors $D^{h}$ and $D^{v}$ such that $D=D^{h}+D^{v}$ and a component $P$ of $D$ dominates $Y$ if and only if $P\subset\operatorname{Supp}D^{h}$. We call such $D^{h}$ and $D^{v}$ the horizontal part and the vertical part of $D$ respectively. We say that $D$ is vertical over $Y$ if $D=D^{v}$. If $D$ is vertical over $Y$, we say that $D$ is exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $f(Q)=P$ but $Q\not\subset\operatorname{Supp}D$.
For a Weil divisor $D$ on a projective normal variety $X$, we denote by $f_{D}$ the rational map associated to the linear system $|D|$. For a linear system $|D|$, we denote by $\operatorname{Mov}|D|$ and $\operatorname{Fix}|D|$ its moving part and fixed part respectively.
@@ 2. On Weil divisors whose section rings are finitely generated
Let $X$ be a projective normal variety and let $D$ be a (non necessarily Cartier) divisor on $X$, such that $R(X,D):=\bigoplus_{m\geq 0}H^{0}(X,mD)$ is finitely generated and is generated in degree one. Let $Y:=\operatorname{Proj}R(X,D)$. Let $f=f_{D}:X\dashrightarrow\mathbb{P}^{N}$ be the rational map associated to the linear system $|D|$. Let $\pi:X^{\prime}\to X$ be a resolution of singularities, such that the rational map $f^{\prime}:X^{\prime}\dashrightarrow Y$ is a morphism. Let $\operatorname{Ex}(\pi)$ be the codimension one part of the exceptional locus of $\pi$. ∎
@@
By Lemma 2.2 and Theorem 2.1, all statements except (1), (2) and (3) are clear.
@@
Since $(K_{X}+\Delta)$ is big, there is an ample $\mathbb{Q}$-divisor AA and an effective $\mathbb{Q}$-divisor $E$ such that $(K_{X}+\Delta)\sim_{\mathbb{Q}}A+E$. If $0<\epsilon\ll 1$, then $(X,\Delta+\epsilon E)$ is klt and $(K_{X}+\Delta+\epsilon E)\sim_{\mathbb{Q}}(1-\epsilon)(A+E)+\epsilon A$ is ample. ∎
@@
Since $K_{X}$ is big and $f^{\prime}$ has connected fibers, we have $f^{\prime}$ is a birational morphism. Hence $F$ and $\operatorname{Ex}(\pi)$ are exceptional over $Y$. Hence the rational map $f:X\dashrightarrow Y$ is a birational contraction, i.e., there are no $f^{-1}$-exceptional divisors on $Y$. By
@@
By [BCHM, Corollary 1.4.3], there is a projective klt pair $(Z,\Delta_{Z})$ and a birational morphism $g:Z\to Y$ such that $K_{Z}+\Delta_{Z}\sim_{\mathbb{Q}}g^{*}K_{Y}$, and the induced rational map $Z\dashrightarrow X$ is small. Since $K_{Y}$ is ample, we have $(K_{Z}+\Delta_{Z})$ is nef and big. Hence $Z$ is of Fano type by Lemma 3.1. By [Birkar2016FanotypeFibraiton, Lemma 2.4], $X$ is of Fano type. ∎

### 16 Mar 2026 v2

@@ 1. Introduction
In [zhuang2024direct], Zhuang proved that pure images of klt type singularities are of klt type. The following klt type criterion is crucial in his proof.
@@ Acknowledgments
The author thanks Professor Chen Jiang for providing a simpler proof of Theorem 1.2. The author acknowledges Professor Mao Sheng for the financial support through the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@
Let $f:X\to Y$ be a projective morphism with connected fibers between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. There are effective $\mathbb{Q}$-divisors $D^{h}$ and $D^{v}$ such that $D=D^{h}+D^{v}$ and a component $P$ of $D$ dominates $Y$ if and only if $P\subset\operatorname{Supp}D^{h}$. We call such $D^{h}$ and $D^{v}$ the horizontal part and the vertical part of $D$, respectively. We say that $D$ is vertical over $Y$ if $D=D^{v}$. If $D$ is vertical over $Y$, we say that $D$ is exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $f(Q)=P$ but $Q\not\subset\operatorname{Supp}D$.
For a Weil divisor $D$ on a projective normal variety $X$, we denote by $f_{D}$ the rational map associated to the linear system $|D|$. For a linear system $|D|$, we denote by $\operatorname{Mov}|D|$ and $\operatorname{Fix}|D|$ its moving part and fixed part, respectively.
@@ 2. On Weil divisors whose section rings are finitely generated
Let $X$ be a projective normal variety and let $D$ be a (non-necessarily Cartier) divisor on $X$, such that $R(X,D):=\bigoplus_{m\geq 0}H^{0}(X,mD)$ is finitely generated and is generated in degree one. Let $Y:=\operatorname{Proj}R(X,D)$. Let $f=f_{D}:X\dashrightarrow\mathbb{P}^{N}$ be the rational map associated to the linear system $|D|$. Let $\pi:X^{\prime}\to X$ be a resolution of singularities, such that the rational map $f^{\prime}:X^{\prime}\dashrightarrow Y$ is a morphism. Let $\operatorname{Ex}(\pi)$ be the codimension one part of the exceptional locus of $\pi$. ∎
@@
By Lemma 2.2 and Theorem 2.1, all statements except (1), (2), and (3) are clear.
@@
Since $(K_{X}+\Delta)$ is big, there are an ample $\mathbb{Q}$-divisor AA and an effective $\mathbb{Q}$-divisor $E$ such that $(K_{X}+\Delta)\sim_{\mathbb{Q}}A+E$. If $0<\epsilon\ll 1$, then $(X,\Delta+\epsilon E)$ is klt and $(K_{X}+\Delta+\epsilon E)\sim_{\mathbb{Q}}(1-\epsilon)(A+E)+\epsilon A$ is ample. ∎
Lemma 3.2.
Let $f:X\to Y$ be a birational morphism between projective normal varieties. If $X$ is of Fano type, then $Y$ is of Fano type.
Proof.
Let $\Delta$ and AA be effective $\mathbb{Q}$-divisors such that AA is ample, $(X,\Delta+A)$ is klt and $K_{X}+\Delta+A\sim_{\mathbb{Q}}0$. By $K_{X}+\Delta+A\sim_{\mathbb{Q}}f^{*}(K_{Y}+f_{*}\Delta+f_{*}A)$, we have $(Y,f_{*}\Delta+f_{*}A)$ is klt. Since $f_{*}A$ is big, there are an ample effective $\mathbb{Q}$-divisor $A^{\prime}$ and an effective $\mathbb{Q}$-divisor $E$ such that $f_{*}A\sim_{\mathbb{Q}}A^{\prime}+E$. If $0<\epsilon\ll 1$, then $(Y,f_{*}\Delta+(1-\epsilon)f_{*}A+\epsilon E)$ is klt and $(K_{Y}+f_{*}\Delta+(1-\epsilon)f_{*}A+\epsilon E)\sim_{\mathbb{Q}}\epsilon A$ is ample. ∎
@@
Since $K_{X}$ is big and $f^{\prime}$ has connected fibers, we have $f^{\prime}$ is a birational morphism. Hence $F$ and $\operatorname{Ex}(\pi)$ are exceptional over $Y$. Hence, the rational map $f:X\dashrightarrow Y$ is a birational contraction, i.e., there are no $f^{-1}$-exceptional divisors on $Y$. By
@@
By [BCHM, Corollary 1.4.3], there is a projective klt pair $(Z,\Delta_{Z})$ and a birational morphism $g:Z\to Y$ such that $K_{Z}+\Delta_{Z}\sim_{\mathbb{Q}}g^{*}K_{Y}$, and the induced rational map $Z\dashrightarrow X$ is small. Since $K_{Y}$ is ample, we have $(K_{Z}+\Delta_{Z})$ is nef and big. Hence $Z$ is of Fano type by Lemma 3.1. By [Birkar2016FanotypeFibraiton, Lemma 2.4], $X$ is of Fano type.
Another Proof of the “if” part (Provided by Chen Jiang). By [BCHM], there is a projective $\mathbb{Q}$-factorial terminal variety $Y$ and a birational morphism $\pi:X^{\prime}\to X$ such that $K_{X^{\prime}}$ is nef over $X$.
We claim that $\pi_{*}:|-mK_{X^{\prime}}|\to|-mK_{X}|$ is one-one for all $m\geq 0$. Let $s$ be a rational function on $X$ such that $\operatorname{div}_{X}(s)-mK_{X}\geq 0$. Then $\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}}$ is a $\pi$-anti-nef divisor with $\pi_{*}(\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}})=\operatorname{div}_{X}(s)-mK_{X}\geq 0$. By the Negativity Lemma, we have $\operatorname{div}_{X^{\prime}}(s)-mK_{X^{\prime}}\geq 0$. Hence, the claim follows.
Since $X^{\prime}$} is $\mathbb{Q}$-Gorenstein, $R(Y,-K_{X^{\prime}})\cong R(X,-K_{X})$ is finitely generated and $\operatorname{Proj}R(X^{\prime},-K_{X^{\prime}})\cong\operatorname{Proj}R(X,-K_{X})$ is klt, we have $X^{\prime}$ is of Fano type by [CasciniGongyo2013anticanonical, Theorem 1.1]. By Lemma 3.2, $X$ is of Fano type. ∎